Project

General

Profile

27-Aug-2020

Agenda:
1. General status updates/Schedule review.
2. Status of DTC to ROC interface.
3. Status of EPICS on DAQ cluster.
4. Status of timing distribution testing.

Rick -- Working on Event Building debugging. Testing one DTC large and others short. Also trying to support Iris and Jason. Daq11 and daq12 heat sinks interfere with MOLEX connectors. Need to confirm with Jason which DTCs got the new MOLEX connectors. The RTF was not getting signal from CFO, Rick realized the CFO sends AC coupled. Greg modified a few of the RTF input circuits. Iris was not able to see any input signals.

Eric -- working on building otsdaq/mu2e on CentOS8 daq16. There are now releases of all mu2e repos, so now you only need to have the repo you are modifying at the test stand. Vivian was having a few issues preparing for moving daq09.

Vivian -- $HOSTNAME was not translating properly when transferring to normal mode. Still unable to configure (slow controls tables giving errors). Trying to get to Evan’s DQM listening on a port, and was not receiving data, but cant get back to configured to check.

Antonio -- Working on XY widget plot (hopefully 1 or 2 days away). Checked in completed work on image map. Found free online tool to make image maps: https://www.image-map.net/

Bertrand --
Craig -- Simon Corrodi plans to resume vertical slice tests.
Gary -- We need to get detectors into production. Should be high priority. Rick is available Tuesday and Wednesday. Craig asking if there is a potential for funding more of Rick’s time in 2020. Tracker Electronics had discussion on sync of front-ends. Vadim wants to look at synchronizing two DRACs.

Glenn -- Working on Rack monitor (FCC temperature monitor), getting distracted by other things (CRV FEBs not working for Ralf).

Greg -- We do not know FY21 walkdowns for Ops. Guidance is flat.
Lei --
Yujing -- working with Vivian.
Ryan -- Following Iris.